# Some remarks on the exponential separation and dimension preserving approximation for sets and measures

### v1

@@ Abstract.
In the dimension theory of sets and measures, a recent breakthrough happened due to Hochman, who introduced the exponential separation condition and proved the Hausdorff dimension result for invariant sets and measures generated by similarities on the real line. Following this groundbreaking work, we make a modest contribution by weakening the condition. We also define some sets in the class of all nonempty compact sets using the Assouad and Hausdorff dimensions and subsets of measures in the Borel probability measures using the $L^{q}$ dimension and prove their density in the respective spaces.
@@ 1. Introduction
  • IFS with bi-Lipschitz mappings $(s_{i}\|x-y\|_{2}\leq\|f_{i}(x)-f_{i}(y)\|_{2}\leq t_{i}\|x-y\|_{2})$ on $\mathbb{R}^{m}$ and SSC, we have $s\leq\dim_{\mathcal{H}}A\leq t,$ where $s$ and $t$ satisfies $\sum_{i=1}^{N}s_{i}^{s}=1$ and $\sum_{i=1}^{N}t_{i}^{t}=1,$ respectively, see [11, Proposition 9.6-9.7]. It is noted that the upper bound holds without any separation condition, but the lower bound does not hold even if we assume OSC. Additionally, it does not serve as a bound for the Assouad dimension. For example, visit [14, Section 8.5].
@@
  • Generally, $\dim_{\mathcal{H}}A\leq\dim_{\mathcal{A}}A,$ but if $A$ is Ahlfors regular, then $\dim_{\mathcal{H}}A=\dim_{\mathcal{A}}A,$ see [14, Theorem 6.4.1]. IFS with similitudes on $\mathbb{R},$ either $\dim_{\mathcal{H}}A=\dim_{\mathcal{A}}A$ or $\dim_{\mathcal{A}}A=1.$
@@
  • In $\mathbb{R}$ with similarities and ESC, Shmerkin computed $\dim_{L^{q}}\mu=\min\{1,\tilde{\tau}(q)/(q-1)\},$ see [30, Theorem 6.6].
  • IFS with similarity and OSC implies $\dim_{\mathcal{H}}\mu=\min\{m,\dim_{S}\mu\}$, see [10]. In this case, under SSC, Fraser and Howroyd computed $\dim_{\mathcal{A}}\mu,$ see [14, Theorem 7.4.1].
It is well known that the Hausdorff dimension decreases under Lipschitz mapping ([11, Corollary 2.4]), but this is not the case with the Assouad dimension ([14, Corollary 7.2.2]). Bedford [6] in his PhD thesis, studied the box dimension of self-similar sets. Recent work in the dimension estimation of self-similar sets can be explored in [35]. Initially, a special class of self-affine sets known as self-affine carpets, such as Bedford-McMullen carpets, is studied; for details, see [15]. Several recent works in the direction of estimation of the Hausdorff dimension of self-affine sets and measures can be found in [4, 12, 19, 21, 31].
@@
  • The weaker version of the Hochman result [18, Theorem 1.4] is shown in Theorem 3.15.
@@ 2. Preliminaries
Let $\mathcal{K}(\mathbb{R}^{m})$ denote the collection of all nonempty compact subsets of $\mathbb{R}^{m}$ in the Euclidean norm $\|.\|_{2}.$ The Hausdorff distance $\mathfrak{H}$ between any $A,B\in\mathcal{K}(\mathbb{R}^{m})$ is elaborated as
@@
Let $\mathcal{P}(\mathbb{R}^{m})$ be the set of probability measures on $\mathbb{R}^{m}$ with the metric defined as
$d_{L}(\mu,\nu)=\sup_{Lip(f)\leq 1}\Big{\{}\Big{|}\int_{\mathbb{R}^{m}}fd\mu-\int_{\mathbb{R}^{m}}fd\nu\Big{|}\Big{\}},$
where $f$ is a Lipschitz function with the Lipschitz constant $Lip(f)$ equal to $1$, see [7, 25]. Let $\mu,\nu\in\mathcal{P}(\mathbb{R}^{m}).$ Then, the convolution $\mu*\nu$ is given as
@@
[30] Let $q\in(1,\infty).$ If $\mu$ is a probability measure on $\mathbb{R}$ with bounded support, then the $L^{q}$ spectrum of $\mu$ is defined as
@@
If $\mu$ is a probability measure on $\mathbb{R}^{m}$ with bounded support, then the $L^{q}$ spectrum of $\mu$ is defined as
@@
For entropy and $L^{q}$ dimension, researchers consider preferring logarithms on base $2,$ for normalizing the value of dimension. Several recent works on $L^{q}$ dimension can be studied in [2, 24, 30].
@@
We now proceed to estimate the $L^{q}$ dimension of some well-known probability measures, to enhance clarity for new readers.
@@
  • Let $c\in\mathbb{R}.$ Then,
    $\tau(c\delta_{x_{0}},q)=\liminf_{k\rightarrow\infty}\frac{-\log(c)^{q}}{k}=0,\text{ and }\dim_{L^{q}}(c\delta_{x_{0}})=0.$
  • Let $x_{1},x_{2},\ldots,x_{n}$ be distinct points in $\mathbb{R}.$ Then,
@@
For $i=i_{1}i_{2}\ldots i_{n}\in\{1,2,\ldots N\}^{n},$ we denote $h_{i}=h_{i_{1}}\circ h_{i_{2}}\circ\cdots\circ h_{i_{n}}.$ An IFS $\mathcal{I}$ is said to satisfy the exponential separation condition (ESC) if there exists a constant $\delta>0$ such that $\Delta_{n}>\delta^{n}$ for all $n\in\mathbb{N},$ see [19].
We would like to highlight that in some of the articles, the previous definition of ESC is mentioned with all arbitrarily large $n\in\mathbb{N}$ such that $\Delta_{n}>\delta^{n},$ while some others say that $\Delta_{n}>\delta^{n},$ for infinitely many $n\in\mathbb{N},$ for example refer to the literature [3, 18, 30]. Observe that if $\Delta_{n}>\delta^{n}$ for all $n\in\mathbb{N},$ then both are true. In the other case, if $\Delta_{n}>\delta^{n},$ for all $n\geq n_{0},$ then choose $\delta_{*}=\min\{\delta,\Delta_{1},\sqrt{\Delta_{2}}\ldots,\sqrt[n_{0}]{\Delta_{n_{0}}}\}.$ This gives $\Delta_{n}>\delta_{*}^{n},$ for all $n\in\mathbb{N}.$ With reference to Theorem 1.5 of [17], for an IFS on $\mathbb{R}$ consisting of all the similarity mappings with the algebraic parameters, $\Delta_{n}$ is either zero or strictly greater than $\delta^{n},$ for some constant $\delta>0,$ establishing the equivalency among the three. We strongly believe that all three definitions are equivalent in general. Further, note that $\Delta_{n}>\delta^{n}$ for all $n\in\mathbb{N}$ is equivalent to $\Delta_{n}>c\delta^{n}$ for all $n\in\mathbb{N},$ for some constant $c\in\mathbb{R}.$ The direct part is straightforward. For the converse, consider two cases. First $c>1,$ then it is obvious. In the other case, assume $0<c\leq 1.$ Choose $\delta_{*}=c\delta,$ then $\Delta_{n}>c\delta^{n}>(c\delta)^{n}.$ An example was presented by Garsia [16] as $\{\mathbb{R}:~h_{1}(x)=\gamma^{-1}x+1,h_{2}(x)=\gamma^{-1}x-1\},$ where $\gamma$ is the real root of the equation $x^{3}-x^{2}-2=0.$ Hochman [18] shows that this IFS satisfies ESC, but not OSC. Meanwhile, super-exponential condensation means $\lim_{n\rightarrow\infty}\frac{1}{n}\log\Delta_{n}=-\infty.$ Bárány and Käenmäki [3, Theorem 2.1] showed the existence of an uncountable family of homogeneous self-similar sets of $\mathbb{R}$ that do not have total overlaps and satisfy super-exponential condensation. Baker [1] also independently showed the existence of such self-similar sets in $\mathbb{R}$ using the theory of continued fractions. Using the technique of Baker, Chen [9] provided some more examples in $\mathbb{R}.$
@@ 3. Main Results
Let $\mathcal{I}_{1}=\{\mathbb{R}^{k};h_{1},h_{2},\ldots,h_{N}\}$ and $\mathcal{I}_{2}=\{\mathbb{R}^{m};g_{1},g_{2},\ldots,g_{N}\}$ be two IFSs such that $h_{i}(x)=cU_{i}x+a_{i}$ and $g_{i}(x)=cV_{i}x+b_{i},$ where $c\in(-1,1),a_{i}\in\mathbb{R}^{k},b_{i}\in\mathbb{R}^{m},U_{i}$ and $V_{i}$ are $k\times k$ and $m\times m$ orthogonal matrices, respectively. If $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ satisfy ESC, then the product IFS $\mathcal{I}_{1}\times\mathcal{I}_{2}$ given as
@@
Since IFSs $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ satisfy ESC, there exist constants $\delta_{1}$ and $\delta_{2}>0,$ respectively such that
@@
where $\delta_{3}=\min\{\delta_{1},\delta_{2}\},$ for all $n\in\mathbb{N},$ assuring that $\mathcal{I}_{1}\times\mathcal{I}_{2}$ satisfies ESC. ∎
@@
The previous theorem plays a considerable role in the context of extending the Hochman [18, Theorem 1.4] groundbreaking result to the product of IFSs. Meanwhile, Bárány and Verma [5] introduced a weak ESC; it remains an open question whether the result holds in that case as well.
@@
Then, IFS $\{\mathbb{R}^{m};h_{1},h_{2},\ldots,h_{N}\}$ is said to satisfy the modified ESC if there exists a $\delta>0$ such that $\Delta_{n}^{*}>\delta^{n},$ for all $n\in\mathbb{N}.$
Note 3.4.
It is noted that several researchers attempted to generalise the concept of ESC introduced by Hochman [17, 18]. In a way, Bárány and Verma [5] defined weak ESC for self-similar IFS, ESC for graph-directed self-similar IFS, ESC for common fixed point system. Above all that, the most fundamental advantage of the modified ESC is that it is defined not only for self-similar IFS but also extends to more general classes of IFS, including self-affine IFS and self-conformal IFS.
@@
The second example can be considered as Sierpiński gasket on $\mathbb{R}^{2}.$ Its IFS is elaborated as $\{\mathbb{R}^{2}:~h_{1},h_{2},h_{3}\},$ where
@@
Choose $i=11\ldots 1$ and $j=11\ldots 12\in\{1,2,3\}^{n}.$ Then, $h_{i}(\triangle)=$ an equilateral triangle $\triangle$ with vertices $(0,0),(\frac{1}{2^{n}},0),(\frac{1}{2^{n+1}},\frac{\sqrt{3}}{2^{n+1}})$ and $h_{j}(\triangle)=$ an equilateral triangle $\triangle$ with vertices $(\frac{1}{2^{n}},0),(\frac{1}{2^{n-1}},0),\\ (\frac{3}{2^{n+1}},\frac{\sqrt{3}}{2^{n+1}}).$ Then, $\mathfrak{H}(h_{i}(\triangle),h_{j}(\triangle))=\frac{1}{2^{n}}$ and $\Delta_{n}^{*}=\frac{1}{2^{n}}.$
@@
Consider an IFS consisting of similarity transformations $(h_{i}(x)=c_{i}U_{i}x+a_{i},i=1,2,\ldots,N)$ on $\mathbb{R}^{m},$ then there exists a constant $k\in\mathbb{R}$ such that $\Delta_{n}^{*}\leq k\Delta_{n}.$
@@
Taking supremum both side
@@
This completes the proof.
@@
Consider a homogeneous IFS consisting of similarity transformations $(h_{i}(x)=cx+d_{i},i=1,2,\ldots,N)$ on $\mathbb{R},$ then the above two definitions (mentioned in equations (2.1) and (3.1)) of ESC coincide, that is, $\Delta_{n}=\Delta_{n}^{*}.$
@@
Taking the infimum on both sides over $i,j\in\{1,2,\ldots,N\}^{n},$ we get $\Delta_{n}^{*}=\Delta_{n},$ concluding the proof.
Note 3.9.
If homogeneous IFS on $\mathbb{R}$ satisfies ESC, there exists a constant $\delta>0$ such that $\Delta_{n}>\delta^{n},$ for all $n\in\mathbb{N}.$ This gives $\Delta_{n}^{*}>\delta^{n},$ for all $n\in\mathbb{N}.$ Thus, the homogeneous IFS satisfies modified ESC.
@@
Taking infimum over $i,j\in\{1,2,\ldots,N\}^{n}$ both sides, we get $\frac{1}{\min\{b,1\}}\Delta_{n}^{*}\geq\frac{1}{k_{n}}\Delta_{n}.$ If the IFS satisfies ESC, there exists a constant $\delta>0$ such that $\Delta_{n}>\delta^{n},$ for all $n\in\mathbb{N}.$ This gives $\Delta_{n}^{*}>\frac{\min\{b,1\}}{k_{n}}\delta^{n}=\min\{b,1\}\Big{(}\min_{i\in\{1,2,\ldots,N\}}|c_{i}|\delta\Big{)}^{n},$ for all $n\in\mathbb{N}.$
In the upcoming theorems, several dense subsets of $\mathcal{K}(\mathbb{R}^{m})$ are explored, which will be helpful in the dimension-preserving approximation of sets.
@@
For each $\alpha\in[0,m],$ the set $A_{\alpha}=\{E\in\mathcal{K}(\mathbb{R}^{m}):\dim_{\mathcal{A}}E=\alpha\}$ is dense in $\mathcal{K}(\mathbb{R}^{m}).$
@@
Let $\epsilon>0$ and $E\in\mathcal{K}(\mathbb{R}^{m}).$ Since $E$ is compact subset of $\mathbb{R}^{m},$ there exist $e_{1},e_{2},\ldots,e_{n}\in E$ such that $E\subset\bigcup_{j=1}^{n}B(e_{j},\epsilon),$ where $B(e_{j},\epsilon)$ is the open set centered at $e_{j}$ and radius $\epsilon.$ Subsequently, consider $F=\{e_{1},e_{2},\dots,e_{n}\}\in\mathcal{K}(\mathbb{R}^{m}).$ In view of definition of Hausdorff metric and Example 2.3, we get
@@
This shows that the assertion is true in the case of $\alpha=0.$ Again, assume $\alpha\in(0,m].$ Since the Assouad dimension satisfies the open set property (see Theorem 2.4), $\dim_{\mathcal{A}}B(e_{j},\epsilon)=m.$ Now using [8, Theorem 1], there exists a set $F_{e_{1}}\subset B(e_{1},\epsilon)$ such that $\dim_{\mathcal{A}}F_{e_{1}}=\alpha.$ Define $F^{\prime}=F\cup F_{e_{1}}.$ In view of the finite stability of the Assouad dimension, we get
@@
For each $\alpha\in[0,m],$ the set $H_{\alpha}=\{E\in\mathcal{K}(\mathbb{R}^{m}):\dim_{\mathcal{H}}E=\alpha\}$ is dense in $\mathcal{K}(\mathbb{R}^{m}).$
@@
The set $\mathcal{D}=\{E\in\mathcal{K}(\mathbb{R}^{m}):\dim_{\mathcal{H}}E\neq\dim_{\mathcal{A}}E\}$ is dense in $\mathcal{K}(\mathbb{R}^{m}).$
@@
In case of $m=1,$ using [14, Theorem 2.1.1], one can clearly show that $\mathcal{D}$ is a non-empty subset of $\mathcal{K}(\mathbb{R}^{m}).$ The Bedford-McMullen carpet (denoted as $\mathcal{B}\subset\mathbb{R}^{2}$) has the Hausdorff dimension strictly less than the Assouad dimension (see [15, Theorem 2.1]) is an example in case of $m=2.$ For a higher dimension $(m>2)$, consider the set $\mathcal{B}\times\{(\underset{m-2\text{ times}}{\underbrace{0,0,\dots,0}})\}=\{(x,y,0,0,\ldots,0)\in\mathbb{R}^{m}:(x,y)\in\mathcal{B}\},$ concluding the set $\mathcal{D}$ is non-empty. Further, assume $E\in\mathcal{K}(\mathbb{R}^{m})$ and $\epsilon>0.$ Then, there exist $e_{1},e_{2},\ldots,e_{n}\in E$ such that $E\subset\cup_{j=1}^{n}B(e_{j},\epsilon).$ Observe that the Hausdorff and the Assouad dimensions are invariant under scaling and translation (see [11, 14]). Using translation and scaling, we get $\mathcal{B}\times\{(0,0,\ldots,0)\}\subset B(e_{1},\epsilon).$ Choose $F=\{e_{1},e_{2},\ldots,e_{n}\}\cup(\mathcal{B}\times\{(0,0,\ldots,0)\}).$ Since the Assouad and the Hausdorff dimensions satisfy finite stability, we get
@@
We now mention the following theorem, which is a conclusion of [18, Theorem 1.4].

### v2

@@ Abstract.
In the dimension theory of sets and measures, a recent breakthrough happened due to Hochman, who introduced the exponential separation condition (ESC) and proved the Hausdorff dimension result for invariant sets and measures generated by similarities on the real line. Following this groundbreaking work, we make a modest contribution by weakening the condition. Further, we define the modified ESC using the convex hull of the attractor and show that for homogeneous self-similar IFS on $\mathbb{R},$ both definitions coincide. We also define some sets in the class of all nonempty compact sets using the Assouad and Hausdorff dimensions and subsets of measures in the space of Borel probability measures on $\mathbb{R}^{m}$ using the $L^{q}$ dimension and the Rajchman property, and prove their density in the respective spaces.
@@ 1. Introduction
  • IFS with bi-Lipschitz mappings $(s_{i}\|x-y\|_{2}\leq\|f_{i}(x)-f_{i}(y)\|_{2}\leq t_{i}\|x-y\|_{2})$ on $\mathbb{R}^{m}$ and SSC, we have $s\leq\dim_{\mathcal{H}}A\leq t,$ where $s$ and $t$ satisfies $\sum_{i=1}^{N}s_{i}^{s}=1$ and $\sum_{i=1}^{N}t_{i}^{t}=1,$ respectively, see [12, Proposition 9.6-9.7]. It is noted that the upper bound holds without any separation condition, but the lower bound does not hold even if we assume OSC. Additionally, it does not serve as a bound for the Assouad dimension. For example, visit [15, Section 8.5].
@@
  • Generally, $\dim_{\mathcal{H}}A\leq\dim_{\mathcal{A}}A,$ but if $A$ is Ahlfors regular, then $\dim_{\mathcal{H}}A=\dim_{\mathcal{A}}A,$ see [15, Theorem 6.4.1]. IFS with similitudes on $\mathbb{R},$ either $\dim_{\mathcal{H}}A=\dim_{\mathcal{A}}A$ or $\dim_{\mathcal{A}}A=1.$
@@
  • In $\mathbb{R}$ with similarities and ESC, Shmerkin computed $\dim_{L^{q}}\mu=\min\{1,\tilde{\tau}(q)/(q-1)\},$ see [38, Theorem 6.6].
  • IFS with similarity and OSC implies $\dim_{\mathcal{H}}\mu=\min\{m,\dim_{S}\mu\}$, see [11]. In this case, under SSC, Fraser and Howroyd computed $\dim_{\mathcal{A}}\mu,$ see [15, Theorem 7.4.1].
It is well known that the Hausdorff dimension decreases under Lipschitz mapping ([12, Corollary 2.4]), but this is not the case with the Assouad dimension ([15, Corollary 7.2.2]). Bedford [7] in his PhD thesis, studied the box dimension of self-similar sets. Recent work in the dimension estimation of self-similar sets can be explored in [43]. Initially, a special class of self-affine sets known as self-affine carpets, such as Bedford-McMullen carpets, is studied; for details, see [16]. Several recent works in the direction of estimation of the Hausdorff dimension of self-affine sets and measures can be found in [5, 13, 20, 24, 39].
In geometric measure theory, the Fourier transform has emerged as a powerful technique for determining the Hausdorff dimension. The interplay between dimension theory and Fourier analysis was first observed by Kaufman [23], who proved one part of Marstrand’s projection theorem using the Fourier transform method. Furthermore, this technique allows researchers to introduce the Fourier dimension $(\dim_{\mathcal{F}})$ of any subset of $\mathbb{R}^{m}.$ For any Borel subset $A$ of $\mathbb{R}^{m},$ we have $\dim_{\mathcal{F}}(A)\leq\dim_{\mathcal{H}}(A)$, and $\dim_{\mathcal{F}}(A)=\dim_{\mathcal{H}}(A)$ iff $A$ is Salem set. Interested readers may refer to the book [25], and the references therein for details. Recent work on the Fourier decay of self-similar measures and homogeneous self-affine measures can be referred to [40, 41].
@@
  • The weaker version of the Hochman result [19, Theorem 1.4] is shown in Theorem 3.14.
@@ 2. Preliminaries
Let $\chi(\mathbb{R}^{m})$ denote the collection of all nonempty compact subsets of $\mathbb{R}^{m}$ in the Euclidean norm $\|.\|_{2}.$ The Hausdorff distance $\mathfrak{H}$ between any $A,B\in\chi(\mathbb{R}^{m})$ is elaborated as
@@
Definition 2.1.
[25] The Fourier transform of any Borel probability measure $\eta$ in $\mathbb{R}^{m}$ is defined as
$\hat{\eta}(\zeta)=\int e^{-2\pi i\zeta.y}d\eta(y),~~~~\zeta\in\mathbb{R}^{m}.$
Recall that the Fourier transform of convolution $\eta*\nu$ has the property $\widehat{\eta*\nu}(\zeta)=\hat{\eta}(\zeta)\hat{\nu}(\zeta),$ for details; see [25]. It is said that $\eta$ is a Rajchman measure if $\lim_{|\zeta|\to\infty}|\hat{\eta}(\zeta)|=0.$
@@
Let $\mathcal{P}(\mathbb{R}^{m})$ be the set of Borel probability measures on $\mathbb{R}^{m}$ with the metric defined as
$d_{L}(\mu,\nu)=\sup_{f\in\text{Lip}_{1}}\Big\{\Big|\int_{\mathbb{R}^{m}}fd\mu-\int_{\mathbb{R}^{m}}fd\nu\Big|\Big\},$
where $\text{Lip}_{1}$ is a set of all real-valued Lipschitz functions on $\mathbb{R}^{m}$ with the Lipschitz constant less than or equal to $1$, see [8, 30]. Let $\mu,\nu\in\mathcal{P}(\mathbb{R}^{m}).$ Then, the convolution $\mu*\nu$ is given as
@@
We now define $L^{q}$ dimension of a Borel probability measure with bounded support. This concept was initially introduced by Rényi [36], and is one of the fundamental ingredients in the study of multifractal formalism; see [22].
@@
[38] Let $q\in(1,\infty).$ If $\mu$ be a Borel probability measure on $\mathbb{R}$ with bounded support, then the $L^{q}$ spectrum of $\mu$ is defined as
@@
If $\mu$ is a Borel probability measure on $\mathbb{R}^{m}$ with bounded support, then the $L^{q}$ spectrum of $\mu$ is defined as
@@
For entropy and $L^{q}$ dimension, researchers consider preferring logarithms on base $2,$ for normalizing the value of dimension. As $q\rightarrow 1^{+},$ the $L^{q}$ dimension tends to the Hausdorff dimension of the measure. Shmerkin [38] computed $L^{q}$ dimension of self-similar measure under ESC. Peres and Solomyak [31] computed the $L^{q}$ dimension of self-conformal measures, without any separation assumptions. Several recent works on $L^{q}$ dimension can be studied in [2, 27, 28, 29].
@@
We now proceed to estimate the $L^{q}$ dimension of some well-known Borel probability measures, to enhance clarity for new readers.
@@
  • Let $x_{1},x_{2},\ldots,x_{n}$ be distinct points in $\mathbb{R}$ and $c_{1},c_{2},\ldots,c_{n}>0.$ Then,
@@
For $i=i_{1}i_{2}\ldots i_{n}\in\{1,2,\ldots N\}^{n},$ we denote $h_{i}=h_{i_{1}}\circ h_{i_{2}}\circ\cdots\circ h_{i_{n}}.$ An IFS $\mathcal{I}$ is said to satisfy the exponential separation condition (ESC) if there exists a constant $\delta>0$ such that $\Delta_{n}>\delta^{n}$ for infinitely many $n\in\mathbb{N},$ where as $\mathcal{I}$ is said to satisfy the strong ESC if there exists a constant $\delta>0$ such that $\Delta_{n}>\delta^{n},$ for all $n\in\mathbb{N},$ see [3, 19, 20, 38].
An example was presented by Garsia [17] as $\{\mathbb{R}:~h_{1}(x)=\gamma^{-1}x+1,h_{2}(x)=\gamma^{-1}x-1\},$ where $\gamma$ is the real root of the equation $x^{3}-x^{2}-2=0.$ Hochman [19] shows that this IFS satisfies ESC, but not OSC. Meanwhile, super-exponential condensation means $\lim_{n\rightarrow\infty}\frac{1}{n}\log\Delta_{n}=-\infty.$ Bárány and Käenmäki [3, Theorem 2.1] showed the existence of an uncountable family of homogeneous self-similar sets of $\mathbb{R}$ that do not have total overlaps and satisfy super-exponential condensation. Baker [1] also independently showed the existence of such self-similar sets in $\mathbb{R}$ using the theory of continued fractions. Using the technique of Baker, Chen [10] provided some more examples in $\mathbb{R}.$
@@ 3. Main Results
Let $\mathcal{I}_{1}=\{\mathbb{R}^{k};h_{1},h_{2},\ldots,h_{N}\}$ and $\mathcal{I}_{2}=\{\mathbb{R}^{m};g_{1},g_{2},\ldots,g_{N}\}$ be two IFSs such that $h_{i}(x)=cU_{i}x+a_{i}$ and $g_{i}(x)=cV_{i}x+b_{i},$ where $c\in(-1,1),a_{i}\in\mathbb{R}^{k},b_{i}\in\mathbb{R}^{m},U_{i}$ and $V_{i}$ are $k\times k$ and $m\times m$ orthogonal matrices, respectively. If $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ satisfy strong ESC, then the product IFS $\mathcal{I}_{1}\times\mathcal{I}_{2}$ given as
@@
Since IFSs $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ satisfy strong ESC, there exist constants $\delta_{1}$ and $\delta_{2}>0,$ respectively such that
@@
where $\delta_{3}=\min\{\delta_{1},\delta_{2}\},$ for all $n\in\mathbb{N},$ assuring that $\mathcal{I}_{1}\times\mathcal{I}_{2}$ satisfies strong ESC. ∎
@@
The previous theorem plays a considerable role in the context of extending the Hochman [19, Theorem 1.4] groundbreaking result to the product of IFSs.
@@
Then, IFS $\{\mathbb{R}^{m};h_{1},h_{2},\ldots,h_{N}\}$ is said to satisfy the modified ESC if there exists a $\delta>0$ such that $\Delta_{n}^{*}>\delta^{n},$ for infinitely many $n\in\mathbb{N}.$ Similarly, we can define modified strong ESC.
It is worth emphasizing that our definition of ESC is natural from a geometric perspective and significantly convenient. Computability and applicability of ESC and modified ESC might be seen as one being more algebraic in nature and the other is more geometric in nature. It is noted that several researchers [4, 6, 35] attempt to generalize and weaken the concept of ESC introduced by Hochman [18, 19]. In a way, Bárány and M. Verma [4] defined weak ESC for self-similar IFS, ESC for graph-directed self-similar IFS, ESC for common fixed point system. Above all that, the most fundamental advantage of the modified ESC is that it is defined not only for self-similar IFS but also extends to more general classes of IFS, including self-affine IFS and self-conformal IFS.
@@
This clearly shows that the IFS satisfies both modified ESC and modified strong ESC. The second example can be considered as Sierpiński gasket on $\mathbb{R}^{2}.$ Its IFS is elaborated as $\{\mathbb{R}^{2}:~h_{1},h_{2},h_{3}\},$ where
@@
Choose $i=11\ldots 1$ and $j=11\ldots 12\in\{1,2,3\}^{n}.$ Then, $h_{i}(\triangle)=$ an equilateral triangle $\triangle$ with vertices $(0,0),(\frac{1}{2^{n}},0),(\frac{1}{2^{n+1}},\frac{\sqrt{3}}{2^{n+1}})$ and $h_{j}(\triangle)=$ an equilateral triangle $\triangle$ with vertices $(\frac{1}{2^{n}},0),(\frac{1}{2^{n-1}},0),\\ (\frac{3}{2^{n+1}},\frac{\sqrt{3}}{2^{n+1}}).$ Then, $\mathfrak{H}(h_{i}(\triangle),h_{j}(\triangle))=\frac{1}{2^{n}}$ and $\Delta_{n}^{*}=\frac{1}{2^{n}}.$ This clearly shows that the IFS satisfies both modified ESC and modified strong ESC.
@@
Consider an IFS consisting of similarity transformations $(h_{i}(x)=c_{i}U_{i}x+a_{i},i=1,2,\ldots,N)$ on $\mathbb{R}^{m},$ then there exists a constant $k_{*}\in\mathbb{R}$ such that $\Delta_{n}^{*}\leq k_{*}\Delta_{n},$ for all $n\in\mathbb{N}.$
@@
Taking the supremum of both sides, we get
@@
Clearly, this holds for all $n\in\mathbb{N}.$ This completes the proof.
@@
Consider a homogeneous IFS consisting of similarity transformations $(h_{i}(x)=cx+d_{i},i=1,2,\ldots,N)$ on $\mathbb{R},$ then the above two definitions (mentioned in equations (2.1) and (3.1)) of ESC coincide, that is, $\Delta_{n}=\Delta_{n}^{*},$ for all $n\in\mathbb{N}.$
@@
Taking the infimum on both sides over $i,j\in\{1,2,\ldots,N\}^{n},$ we get $\Delta_{n}^{*}=\Delta_{n},$ for all $n\in\mathbb{N},$ concluding the proof.
Note 3.8.
If homogeneous IFS on $\mathbb{R}$ satisfies ESC, there exists a constant $\delta>0$ such that $\Delta_{n}>\delta^{n},$ for infinitely many $n\in\mathbb{N}.$ This gives $\Delta_{n}^{*}>\delta^{n},$ for infinitely many $n\in\mathbb{N}.$ Thus, the homogeneous IFS satisfies modified ESC.
@@
Taking infimum over $i,j\in\{1,2,\ldots,N\}^{n}$ both sides, we get $\frac{1}{\min\{b,1\}}\Delta_{n}^{*}\geq\frac{1}{k_{n}}\Delta_{n}.$ If the IFS satisfies ESC, there exists a constant $\delta>0$ such that $\Delta_{n}>\delta^{n},$ for infinitely many $n\in\mathbb{N}.$ This gives $\Delta_{n}^{*}>\frac{\min\{b,1\}}{k_{n}}\delta^{n}=\min\{b,1\}\Big(\min_{i\in\{1,2,\ldots,N\}}|c_{i}|\delta\Big)^{n},$ for infinitely many $n\in\mathbb{N},$ claiming that the IFS satisfies modified ESC.
In the upcoming theorems, several dense subsets of $\chi(\mathbb{R}^{m})$ are explored, which will be helpful in the dimension-preserving approximation of sets.
@@
For each $\alpha\in[0,m],$ the set $A_{\alpha}=\{E\in\chi(\mathbb{R}^{m}):\dim_{\mathcal{A}}E=\alpha\}$ is dense in $\chi(\mathbb{R}^{m}).$
@@
Let $\epsilon>0$ and $E\in\chi(\mathbb{R}^{m}).$ Since $E$ is compact subset of $\mathbb{R}^{m},$ there exist $e_{1},e_{2},\ldots,e_{n}\in E$ such that $E\subset\bigcup_{j=1}^{n}B(e_{j},\epsilon),$ where $B(e_{j},\epsilon)$ is the open set centered at $e_{j}$ and radius $\epsilon.$ Subsequently, consider $F=\{e_{1},e_{2},\dots,e_{n}\}\in\chi(\mathbb{R}^{m}).$ In view of definition of Hausdorff metric and Example 2.4, we get
@@
This shows that the assertion is true in the case of $\alpha=0.$ Again, assume $\alpha\in(0,m].$ Since the Assouad dimension satisfies the open set property (see Theorem 2.5), $\dim_{\mathcal{A}}B(e_{j},\epsilon)=m.$ Now using [9, Theorem 1], there exists a set $F_{e_{1}}\subset B(e_{1},\epsilon)$ such that $\dim_{\mathcal{A}}F_{e_{1}}=\alpha.$ Define $F^{\prime}=F\cup F_{e_{1}}.$ In view of the finite stability of the Assouad dimension, we get
@@
For each $\alpha\in[0,m],$ the set $H_{\alpha}=\{E\in\chi(\mathbb{R}^{m}):\dim_{\mathcal{H}}E=\alpha\}$ is dense in $\chi(\mathbb{R}^{m}).$
@@
The set $\mathcal{D}=\{E\in\chi(\mathbb{R}^{m}):\dim_{\mathcal{H}}E\neq\dim_{\mathcal{A}}E\}$ is dense in $\chi(\mathbb{R}^{m}).$
@@
In case of $m=1,$ using [15, Theorem 2.1.1], one can clearly show that $\mathcal{D}$ is a non-empty subset of $\chi(\mathbb{R}^{m}).$ The Bedford-McMullen carpet (denoted as $\mathcal{B}\subset\mathbb{R}^{2}$) has the Hausdorff dimension strictly less than the Assouad dimension (see [16, Theorem 2.1]) is an example in case of $m=2.$ For a higher dimension $(m>2)$, consider the set $\mathcal{B}\times\{(\underset{m-2\text{ times}}{\underbrace{0,0,\dots,0}})\}=\{(x,y,0,0,\ldots,0)\in\mathbb{R}^{m}:(x,y)\in\mathcal{B}\},$ concluding the set $\mathcal{D}$ is non-empty. Further, assume $E\in\chi(\mathbb{R}^{m})$ and $\epsilon>0.$ Then, there exist $e_{1},e_{2},\ldots,e_{n}\in E$ such that $E\subset\cup_{j=1}^{n}B(e_{j},\epsilon).$ Observe that the Hausdorff and the Assouad dimensions are invariant under scaling and translation (see [12, 15]). Using translation and scaling, we get $\mathcal{B}\times\{(0,0,\ldots,0)\}\subset B(e_{1},\epsilon).$ Choose $F=\{e_{1},e_{2},\ldots,e_{n}\}\cup(\mathcal{B}\times\{(0,0,\ldots,0)\}).$ Since the Assouad and the Hausdorff dimensions satisfy finite stability, we get
@@
We now mention the following theorem, which is a conclusion of [19, Theorem 1.4].